1) A swimming bath can be filled by two pipes,A and B, in 15 hours. If A only is used, it takes 24 hours to fill the bath. How long will it take if B only is used.

2) A liquid X contains 38% alcohol and a liquid Y contains 71% alcohol. In what ratio must X and Y be mixed to obtain a liquid containing 50% alcohol?


1
Expert's answer
2021-06-24T18:28:16-0400

1)

Let "x=" the time taken to fill a swimming bath by pipe B.

Then


"(\dfrac{1}{24}+\dfrac{1}{x})\cdot15=1"

"\dfrac{1}{x}=\dfrac{1}{15}-\dfrac{1}{24}"


"\dfrac{1}{x}=\dfrac{1}{40}"

"x=40"

If B only is used, it takes 40 hours to fill the bath. 


2)

Let "x=" the amount of a liquid X in litres,  "y=" the amount of a liquid Y in litres.

Then


"0.38x+0.71y=0.5(x+y)"

"0.12x=0.21y"

"4x=7y"

A liquid X and a liquid Y must be mixed in ratio "7\div 4" to obtain a liquid containing 50% alcohol.